A 100 amp 120V load is 18 m from the hydro service. It has been decided to feed the system with #2AWG copper RWU90 cable. Determine the maximum length of run between the load and the distribution centre for the 120 V load. A 2% voltage drop is acceptable. What distance can the #2AWG cable be run? Is the #2AWG cable sufficient for the job? question from hydrocodes and standards
Answer #1

ANSWER:

point to remeber:

1. total run of cable between source and load = 2* one way distance between source and load

2. Voltage drop of cable

= Max. load ampere on cable * resistance(for copper/aluminium conductor) of cable (in ohm/ km) * total run of cable(in km)

3.voltage drop % = (voltage drop/source volatge)* 100 %

Voltage drop % = 1.85/120 = 0.0154 = 1.540%. So, #2AWG Cable is sufficient for the job. And, this cable can be used for total